# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 3.5. Open Kulikov degenerations
Recall from Definition 3.19 the notion of a “generic” open Kulikov surface. We define below open Kulikov degenerations as “semistable smoothings” of a generic open Kulikov surfaces.
@@
An open Kulikov degeneration is a proper morphism $\pi:\mathcal{X}\rightarrow\Delta:=\mathrm{Spf}\,\mathbb{C}[\![t]\!]$ of formal schemes such that,
@@
    the total space $\mathcal{X}$ is smooth,
@@
    the central fiber $\mathcal{X}_{0}$ is a reduced normal crossing divisor in $\mathcal{X}$,
@@
    the central fiber $\mathcal{X}_{0}$ is a generic open Kulikov surface.
@@
By Lemma 3.10, we have $H^{1}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so it follows from [95, Theorem 1.4] (see also [26, Theorem 3.1]) that $f$ is a flat deformation of $f_{0}$. To prove that $\overline{\mathcal{X}}$ has Gorenstein canonical singularities, first note that $\mathcal{X}$ has trivial canonical line bundle – this can be shown analogously as in [33, Theorem 5.10, 3)], shown in the context of degenerations of K3 surfaces. Hence, since $\mathcal{X}$ is a smooth threefold with trivial canonical class, by [59, Corollary 1.5] it follows that $\overline{\mathcal{X}}$ has Gorenstein canonical singularities. Note that [59, Corollary 1.5] is stated in the context of projective morphism between normal algebraic varieties. In this context, the result follows from the base point free theorem [59, Theorem 1.3]. The same argument shows that the result for projective morphisms of formal schemes follows from the corresponding version of the base point free theorem in [74, §11].
It remains to show that $\overline{\mathcal{X}}\rightarrow\Delta$ is a smoothing of $\mathcal{X}_{0}$. It is enough to show that the rigid-analytic generic fiber $\overline{\mathcal{X}}_{\eta}$ of $\overline{\mathcal{X}}$ is smooth. Denote by $\mathcal{X}_{\eta}$ the rigid-analytic generic fiber of $\mathcal{X}$, which is smooth since $\mathcal{X}\rightarrow\Delta$ is smoothing. Standard results on birational geometry of surfaces generalize to rigid-analytic surfaces by [93, §5]. Hence, it follows from the triviality of the canonical class of $\mathcal{X}_{\eta}$ has trivial that the birational map $\mathcal{X}_{\eta}\rightarrow\overline{\mathcal{X}}_{\eta}$ can only contract $(-2)$-curves. Hence, it is enough to show that $\mathcal{X}_{\eta}$ does not contain any $(-2)$-curve. If $C_{\eta}$ were a $(-2)$-curve in $\mathcal{X}_{\eta}$, then up to replacing $\mathcal{X}$ by a log resolution $\widetilde{\mathcal{X}}$ of a base change, $C_{\eta}$ would admit a model with central fiber a curve $C_{0}$ in $\widetilde{\mathcal{X}}_{0}$ containing no 0-dimensional stratum of $\widetilde{\mathcal{X}}_{0}$. Such a curve $C_{0}$ would necessarily be either an internal $(-2)$-curve or a degenerate $(-2)$-curve as in Definition 3.17, in contradiction with the assumption that the open Kulikov surface $\mathcal{X}_{0}$ is generic as in Definition 3.19. ∎
@@
Since $\overline{\mathcal{X}}$ is an affine formal smoothing, it follows from Elkik’s theorem [29, Theorem 7, p582] – see also [91, Proposition 3.3.1], that $\overline{\mathcal{X}}$ is algebraizable over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$. Moreover, $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a formal blow-up, and so is algebraizable in the category of algebraic spaces over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ by [19, Theorem 3.2]. Since $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is projective, it follows that $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is in fact algebraizable in the category of schemes over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$.
@@
The only non-immediate implication is i) implies ii). By Lemma 3.10, we have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so there are no obstructions to formal deformations over $\Delta$ of an ample line bundle on $\mathcal{X}_{0}$, and so the result follows. ∎
@@
    There exist positive integers $b_{j}$ such that
    $(\sum_{j}b_{j}D_{j})\cdot D_{i}>0$
    for all $i$.
@@
The equivalence between (i), (ii), (iii) and (iv) is the equivalence between (1.1), (1.2), (1.3) and (1.4) in [40, Lemma 6.9].
If $W=-(K_{Y}+\sum_{i}c_{i}D_{i})$ is big and nef, then $-K_{Y}=W+\sum_{i}c_{i}D_{i}$ is the sum of a big divisor with an effective divisor, and so $-K_{Y}$ big. This shows that (iv) implies (v). Finally, we show that (v) implies (iii). If $-K_{Y}$ is big, then $\sum_{i}D_{i}=-K_{Y}$ is big, and so the result follows from the proof that (1.1) implies (1.2) in the proof of [40, Lemma 6.9]. ∎
@@ 3.6. Mori fans of quasi-projective open Kulikov degenerations
In this section, we show that the total space of a quasi-projective open Kulikov degeneration is a (relative) Mori dream space as in [52] for the absolute case, and [12, 79] for the relative case. It follows that one can attach to any quasi-projective open Kulikov degeneration the corresponding Mori fan.
@@
In what follows we describe the Mori fan associated to an open Kulikov degeneration. Note that for any Mori dream space $V$ over $W$, where $V$ and $W$ are normal algebraic varieties with a projective morphism $f:V\to W$ there is an associated fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$, known as the Mori fan, with support the cone of $f$-effective divisors, $\mathrm{Eff}(V/W)$ [52, 79]. The maximal dimensional cones of the Mori fan are in one-to-one correspondence with the birational contractions $g:V\dashrightarrow V^{\prime}$ over $W$ where $V^{\prime}$ is $\mathbb{Q}$-factorial. For such a birational contraction, the corresponding maximal cone of the Mori fan is given by
$g^{*}\mathrm{Nef}(V^{\prime}/W)+\mathrm{Ex}(g)$
where by $\mathrm{Ex}(g)$ we denote the cone in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ spanned by all $g$-exceptional divisors. Note that, for a birational projective morphism $f:V\to W$, we have $\mathrm{Eff}(V/W)=\mathrm{Pic}(V/W)_{\mathbb{R}}$. Thus, the Mori fan is a complete fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$.
Working in the set up of formal Mori theory as in [74], analogously one has the notion of a complete Mori fan associated to a birational projective morphism $f:V\to W$ such that $V$ is a Mori dream space over $W$. Given an open Kulikov degeneration, with affinization $f\colon\mathcal{X}\to\overline{\mathcal{X}}$, $\mathcal{X}$ is a Mori dream space over $\overline{\mathcal{X}}$ by Proposition 3.36, and so we have an associated Mori fan, which we denote by $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$ in the following sections.
@@ 4.1. Deformation theory of log Calabi–Yau surfaces
The result is well-known if $D$ is empty, see for example [81, Theorem 4.10], [75, Lemma 1], or [94, Proposition 6.4]. On the other hand, if $D$ is non-empty but $X$ is smooth, a proof can be found in [34, Proposition 3.5]. Our proof below of the general case is based on a combination of the techniques used for these two known special cases.
Let $U$ be the non-empty open subset of $X$ where both $X$ and $D$ are smooth, and denote by $j:U\hookrightarrow X$ the inclusion. By assumption, the complement $X\setminus D$ consists of isolated points in the surface $X$, and so if of codimension two. Consider the complex
$A:=(j_{\star}T_{X}|_{U}\longrightarrow j_{\star}N_{D|U})\,$
where $T_{X}|_{U}$ is the restriction to $U$ of the tangent sheaf $T_{X}$, placed in degree 00, and $N_{D|U}=\mathcal{O}_{D\cap U}(D\cap U)$ is the normal sheaf to $D\cap U$ in $D\cap U$, placed in degree $1$. Using as in the proof of [81, Theorem 4.10] a Čech cover of $X$ by affine subsets centered on the isolated singularities and with double and triple intersections contained in $U$, one shows that if the hypercohomology group $\mathbb{H}^{2}(X,A)$ vanishes, then there are no local-to-global obstructions to deformations of $(X,D)$. When $X$ is smooth, details of this argument can be found in [87, Proposition 8].
@@
defining $D\cap U$ in $U$. As $T_{X}|_{U}$ is locally free, tensoring by $T_{X}|_{U}$ induces a short exact sequence
$0\rightarrow T_{X}|_{U}(-D\cap U)\rightarrow T_{X}|_{U}\rightarrow T_{X}|_{D\cap U}\rightarrow 0\,.$
Applying $j_{\star}$, we obtain an exact sequence of the form
@@
Consider the complex
$B:=(L\longrightarrow j_{\star}N_{D|U})\,,$
obtained by restricting to $L$ the natural morphism $(j_{\star}T_{X}|_{D\cap U}\longrightarrow j_{\star}N_{D|U})$. Restriction to $D$ induces a map of complexes $A\rightarrow B$. By construction, we have an exact sequence of complexes
@@
The kernel $\mathcal{H}^{0}(B)$ of $B$ is supported in dimension one, so we have
$H^{2}(X,\mathcal{H}^{0}(B))=0\,,$
and the cokernel $\mathcal{H}^{1}(B)$ of $B$ is supported in $X\setminus U$, that is in dimension zero, and so we have also
$H^{1}(X,\mathcal{H}^{1}(B))=0\,.$
It follows from the hypercohomology spectral sequence that $\mathbb{H}^{2}(X,B)=0$, and this concludes the proof. ∎
@@
The following argument is inspired by the proof of [75, Theorem 21]. By Lemma 4.1, it is enough to show that $H^{2}(X,(T_{X}(K_{X}))^{\vee\vee})=0$. A normal surface being Cohen-Macaulay, the reflexive canonical sheaf $\mathcal{O}_{X}(K_{X})$ is the dualizing sheaf of $X$, and so by Serre duality we have
@@
Let $j:U\hookrightarrow X$ be the open immersion of the smooth locus of $X$. As $X$ is normal and $X\setminus U$ is of codimension two, the functors $j_{\star}$ and $j^{\star}$ are inverse equivalences between the category of coherent reflexive sheaves on $X$ and the coherent reflexive sheaves on $U$ [88, Lemma 0EBJ]. As both $(T_{X}(K_{X}))^{\vee\vee}$ and $\mathcal{O}_{X}(K_{X})$ are reflexive, it follows that
@@
Finally, as $X$ has quotient singularities, it follows from Steenbrink Hodge theory for orbifolds [89, §1] that $\dim H^{0}(X,\Omega_{X}^{\vee\vee})=\dim H^{1}(X,\mathcal{O}_{X})$. Hence, we have proved that
@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
Throughout this section we denote by $M\cong\mathbb{Z}^{2}$ the character lattice for $\overline{Y}$, so that $\overline{P}\subseteq M_{\mathbb{R}}$, where $M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ denotes the real vector space associated to $M$. We will first construct a Mumford degeneration associated to the toric polytope $\overline{P}$ endowed with the polyhedral decomposition $\overline{\mathscr{P}}$. As in [77, §3], the union of cones
@@
over each cell $\sigma\in\overline{\mathscr{P}}$ defines a toric fan with support in $M_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$, which we denote by $C(\overline{P})$. We denote by $\mathcal{X}_{\overline{\mathscr{P}}}$ the 3-dimensional toric variety with fan $C(\overline{P})$. By Definition 2.13 of a good polyhedral decomposition, $\overline{\mathscr{P}}$ is a regular decomposition, and so $\mathcal{X}_{\overline{\mathscr{P}}}$ is a quasi-projective toric variety. The projection onto the second factor defines a map of fans from $C(\overline{P})$ to the toric fan of $\mathbb{A}^{1}$, hence induces a toric morphism
$\pi_{\overline{\mathscr{P}}}\colon\mathcal{X}_{\overline{\mathscr{P}}}\to\mathbb{A}^{1}\,.$ (4.3)
The central fiber $\mathcal{X}_{\overline{\mathscr{P}},0}\coloneqq\pi^{-1}(0)$ and the general fiber $\pi^{-1}(t)$ for $t\neq 0$ are described by examining pre-images of the cones $\mathbb{R}_{\geq 0}$ and $\{0\}$ respectively, in the fan of $\mathbb{A}^{1}$ – for details see [77, §3]. Since in our situation all cells $\sigma\subset\overline{\mathscr{P}}$ are bounded, the so called asymptotic cone in [77] consists of a single point, the origin in $M_{\mathbb{R}}\times\mathbb{R}$ and hence the general fiber is the algebraic torus (*$(\mathbb{C}^{*})^{2}$. The central fiber $\mathcal{X}_{\overline{\mathscr{P}},0}$ is a union of irreducible components $X_{\overline{\mathscr{P}}}^{v}$, which are toric surfaces corresponding to vertices $v\in\overline{\mathscr{P}}^{[0]}$,
@@
The associated toric fan for each such a component $\mathcal{X}_{\overline{\mathscr{P}},0}$ corresponding to a vertex $v\in\overline{\mathscr{P}}^{[0]}$ is given by the star of $v$. We denote by $\partial X_{\overline{\mathscr{P}}}^{v}$ the union of intersections of $X_{\overline{\mathscr{P}}}^{v}$ with the other irreducible components of $\mathcal{X}_{\overline{\mathscr{P}},0}$. For every vertex $v\in\overline{\mathscr{P}}^{[0]}$, $\partial X_{\overline{\mathscr{P}}}^{v}$ is the union of toric divisors of $X_{\overline{\mathscr{P}}}^{v}$.

### 30 Mar 2024 v2

@@ 3.5. Open Kulikov degenerations
We define below open Kulikov degenerations as “semistable smoothings” of a generic open Kulikov surfaces.
@@
An open Kulikov degeneration is a proper morphism $\pi:\mathcal{X}\rightarrow\Delta:=\mathrm{Spf}\,\mathbb{C}[\![t]\!]$ of formal schemes satisfying the following conditions.
@@
    The total space $\mathcal{X}$ is smooth.
@@
    The central fiber $\mathcal{X}_{0}$ is a reduced normal crossing divisor in $\mathcal{X}$.
@@
    The central fiber $\mathcal{X}_{0}$ is a generic open Kulikov surface – as in Definition 3.19.
@@
By Lemma 3.10, we have $H^{1}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so it follows from [96, Theorem 1.4] (see also [26, Theorem 3.1]) that $f$ is a flat deformation of $f_{0}$. To prove that $\overline{\mathcal{X}}$ has Gorenstein canonical singularities, first note that $\mathcal{X}$ has trivial canonical line bundle – this can be shown analogously as in [33, Theorem 5.10, 3)], where a similar result is proved in the context of degenerations of K3 surfaces. Hence, since $\mathcal{X}$ is a smooth threefold with trivial canonical class, by [60, Corollary 1.5] it follows that $\overline{\mathcal{X}}$ has Gorenstein canonical singularities. Note that [60, Corollary 1.5] is stated in the context of projective morphism between normal algebraic varieties. In this context, the result follows from the base point free theorem [60, Theorem 1.3]. The same argument shows that the result for projective morphisms of formal schemes follows from the corresponding version of the base point free theorem in [75, §11].
It remains to show that $\overline{\mathcal{X}}\rightarrow\Delta$ is a smoothing of $\mathcal{X}_{0}$. It is enough to show that the rigid-analytic generic fiber $\overline{\mathcal{X}}_{\eta}$ of $\overline{\mathcal{X}}$ is smooth. Denote by $\mathcal{X}_{\eta}$ the rigid-analytic generic fiber of $\mathcal{X}$, which is smooth since $\mathcal{X}\rightarrow\Delta$ is smoothing. Standard results on birational geometry of surfaces generalize to rigid-analytic surfaces by [95, §5]. Hence, it follows from the triviality of the canonical class of $\mathcal{X}_{\eta}$ has trivial that the birational map $\mathcal{X}_{\eta}\rightarrow\overline{\mathcal{X}}_{\eta}$ can only contract $(-2)$-curves. Hence, it is enough to show that $\mathcal{X}_{\eta}$ does not contain any $(-2)$-curve. If $C_{\eta}$ were a $(-2)$-curve in $\mathcal{X}_{\eta}$, then up to replacing $\mathcal{X}$ by a log resolution $\widetilde{\mathcal{X}}$ of a base change, $C_{\eta}$ would admit a model with central fiber a curve $C_{0}$ in $\widetilde{\mathcal{X}}_{0}$ containing no 0-dimensional stratum of $\widetilde{\mathcal{X}}_{0}$. Such a curve $C_{0}$ would necessarily be either an internal $(-2)$-curve or a degenerate $(-2)$-curve as in Definition 3.17, in contradiction with the assumption that the open Kulikov surface $\mathcal{X}_{0}$ is generic as in Definition 3.19. ∎
@@
Since $\overline{\mathcal{X}}$ is an affine formal smoothing, it follows from Elkik’s theorem [29, Theorem 7, p582] – see also [92, Proposition 3.3.1], that $\overline{\mathcal{X}}$ is algebraizable over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$. Moreover, $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a formal blow-up, and so is algebraizable in the category of algebraic spaces over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$ by [19, Theorem 3.2]. Since $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is projective, it follows that $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is in fact algebraizable in the category of schemes over $\mathrm{Spec}\,\mathbb{C}[\![t]\!]$.
@@
The only non-immediate implication is i) implies ii). By Lemma 3.10, we have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, so there are no obstructions to formal deformations over $\Delta$ of an ample line bundle on $\mathcal{X}_{0}$, and so the result follows. ∎
@@
    There exist positive integers $b_{j}$ such that $(\sum_{j}b_{j}D_{j})\cdot D_{i}>0$ for all $i$.
@@
The equivalence between (i), (ii), (iii) and (iv) is the equivalence between (1.1), (1.2), (1.3) and (1.4) in [41, Lemma 6.9].
If $W=-(K_{Y}+\sum_{i}c_{i}D_{i})$ is big and nef, then $-K_{Y}=W+\sum_{i}c_{i}D_{i}$ is the sum of a big divisor with an effective divisor, and so $-K_{Y}$ big. This shows that (iv) implies (v). Finally, we show that (v) implies (iii). If $-K_{Y}$ is big, then $\sum_{i}D_{i}=-K_{Y}$ is big, and so the result follows from the proof that (1.1) implies (1.2) in the proof of [41, Lemma 6.9]. ∎
@@ 3.6. Mori fans of quasi-projective open Kulikov degenerations
In this section, we show that the total space of a quasi-projective open Kulikov degeneration is a (relative) Mori dream space, as in [53] for the absolute case and [12, 80] for the relative case. Consequently, we attach to any quasi-projective open Kulikov degeneration the corresponding Mori fan.
@@
In what follows we describe the Mori fan associated to an open Kulikov degeneration. For any Mori dream space $V$ over $W$, where $V$ and $W$ are normal algebraic varieties with a projective morphism $f:V\to W$ there is an associated fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$, known as the Mori fan, with support the cone of $f$-effective divisors, $\mathrm{Eff}(V/W)$ [53, 80]. Using formal Mori theory as in [75], one can similarly construct a Mori fan for a Mori dream space $V$ over $W$, where $V$ and $W$ are formal schemes and $f:V\rightarrow W$ is a projective morphism. The maximal dimensional cones of the Mori fan are in one-to-one correspondence with the birational contractions $g:V\dashrightarrow V^{\prime}$ over $W$ where $V^{\prime}$ is $\mathbb{Q}$-factorial. For such a birational contraction, the corresponding maximal cone of the Mori fan is given by $g^{*}\mathrm{Nef}(V^{\prime}/W)+\mathrm{Ex}(g)$, where $\mathrm{Ex}(g)$ is the cone in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ spanned by all $g$-exceptional divisors. For a birational projective morphism $f:V\to W$, we have $\mathrm{Eff}(V/W)=\mathrm{Pic}(V/W)_{\mathbb{R}}$, and so the Mori fan is a complete fan in $\mathrm{Pic}(V/W)_{\mathbb{R}}$ in this case. Given an open Kulikov degeneration, with affinization $f\colon\mathcal{X}\to\overline{\mathcal{X}}$, $\mathcal{X}$ is a Mori dream space over $\overline{\mathcal{X}}$ by Proposition 3.36, and so we have an associated complete Mori fan in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})$, denoted by $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$ in the following sections.
@@ 4.1. Deformation theory of log Calabi–Yau surfaces
The result is well-known if $D$ is empty, see for example [82, Theorem 4.10], [76, Lemma 1], or [97, Proposition 6.4]. On the other hand, if $D$ is non-empty but $X$ is smooth, a proof can be found in [34, Proposition 3.5]. Our proof below of the general case is based on a combination of the techniques used for these two known special cases.
Let $U$ be the non-empty open subset of $X$ where both $X$ and $D$ are smooth, and denote by $j:U\hookrightarrow X$ the inclusion. By assumption, the complement $X\setminus D$ consists of isolated points in the surface $X$, and so if of codimension two. Consider the complex $A:=(j_{\star}T_{X}|_{U}\rightarrow j_{\star}N_{D|U})$, where $T_{X}|_{U}$ is the restriction to $U$ of the tangent sheaf $T_{X}$, placed in degree 00, and $N_{D|U}=\mathcal{O}_{D\cap U}(D\cap U)$ is the normal sheaf to $D\cap U$ in $D\cap U$, placed in degree $1$. Using as in the proof of [82, Theorem 4.10] a Čech cover of $X$ by affine subsets centered on the isolated singularities and with double and triple intersections contained in $U$, we deduce that if the hypercohomology group $\mathbb{H}^{2}(X,A)$ vanishes, then there are no local-to-global obstructions to deformations of $(X,D)$. When $X$ is smooth, details of this argument can be found in [88, Proposition 8].
@@
defining $D\cap U$ in $U$. As $T_{X}|_{U}$ is locally free, tensoring by $T_{X}|_{U}$, and then applying $j_{\star}$, we obtain an exact sequence of the form
@@
Consider the complex $B:=(L\longrightarrow j_{\star}N_{D|U})$ obtained by restricting to $L$ the natural morphism $(j_{\star}T_{X}|_{D\cap U}\longrightarrow j_{\star}N_{D|U})$. Restriction to $D$ induces a map of complexes $A\rightarrow B$. By construction, we have an exact sequence of complexes
@@
The kernel $\mathcal{H}^{0}(B)$ of $B$ is supported in dimension one, so we have $H^{2}(X,\mathcal{H}^{0}(B))=0$, and the cokernel $\mathcal{H}^{1}(B)$ of $B$ is supported in $X\setminus U$, that is in dimension zero, and so we have also $H^{1}(X,\mathcal{H}^{1}(B))=0$. It follows from the hypercohomology spectral sequence that $\mathbb{H}^{2}(X,B)=0$, and this concludes the proof. ∎
@@
The following argument is inspired by the proof of [76, Theorem 21]. By Lemma 4.1, it is enough to show that $H^{2}(X,(T_{X}(K_{X}))^{\vee\vee})=0$. A normal surface being Cohen-Macaulay, the reflexive canonical sheaf $\mathcal{O}_{X}(K_{X})$ is the dualizing sheaf of $X$, and so by Serre duality we have
@@
Let $j:U\hookrightarrow X$ be the open immersion of the smooth locus of $X$. As $X$ is normal and $X\setminus U$ is of codimension two, the functors $j_{\star}$ and $j^{\star}$ are inverse equivalences between the category of coherent reflexive sheaves on $X$ and the coherent reflexive sheaves on $U$ [89, Lemma 0EBJ]. As both $(T_{X}(K_{X}))^{\vee\vee}$ and $\mathcal{O}_{X}(K_{X})$ are reflexive, it follows that
@@
Finally, as $X$ has quotient singularities, it follows from Steenbrink Hodge theory for orbifolds [90, §1] that $\dim H^{0}(X,\Omega_{X}^{\vee\vee})=\dim H^{1}(X,\mathcal{O}_{X})$. Hence, it follows that
@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
Throughout this section we denote by $M\cong\mathbb{Z}^{2}$ the character lattice for $\overline{Y}$, so that $\overline{P}\subseteq M_{\mathbb{R}}$, where $M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ denotes the real vector space associated to $M$. We first consider the Mumford toric degeneration $\pi_{\overline{\mathscr{P}}}:\mathcal{X}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ associated to the toric polytope $\overline{P}$ endowed with the polyhedral decomposition $\overline{\mathscr{P}}$. As in [78, §3], $\mathcal{X}_{\overline{\mathscr{P}}}$ is the 3-dimensional toric variety with fan $C(\overline{\mathscr{P}})$ in $M_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$ consisting of the cones
@@
over the cells $\sigma\in\overline{\mathscr{P}}$. Moreover, the toric morphism $\pi_{\overline{\mathscr{P}}}\colon\mathcal{X}_{\overline{\mathscr{P}}}\to\mathbb{A}^{1}$ is induced by the projection $M_{\mathbb{R}}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ onto the second factor. By Definition 2.13 of a good polyhedral decomposition, $\overline{\mathscr{P}}$ is a regular decomposition, and so $\mathcal{X}_{\overline{\mathscr{P}}}$ is a quasi-projective toric variety.
Since all cells $\sigma\in\overline{\mathscr{P}}$ are bounded, the so called asymptotic cone in [78] consists of a single point, the origin in $M_{\mathbb{R}}\times\mathbb{R}$, and hence the fibers $\pi_{\overline{\mathscr{P}}}^{-1}(t)$ for $t\neq 0$ are isomorphic to the algebraic torus (*$(\mathbb{C}^{*})^{2}$. The central fiber $\mathcal{X}_{\overline{\mathscr{P}},0}\coloneqq\pi_{\overline{\mathscr{P}}}^{-1}(0)$ is a union of irreducible components $X_{\overline{\mathscr{P}}}^{v}$, indexed by the vertices $v\in\overline{\mathscr{P}}^{[0]}$, which are toric surfaces with fan the star of $v$ in $\overline{\mathscr{P}}$:
@@
For every vertex $v\in\overline{\mathscr{P}}^{[0]}$, we denote by $\partial X_{\overline{\mathscr{P}}}^{v}$ the union of intersections of $X_{\overline{\mathscr{P}}}^{v}$ with the other irreducible components of $\mathcal{X}_{\overline{\mathscr{P}},0}$, which is also the union of toric divisors of $X_{\overline{\mathscr{P}}}^{v}$.